# MetaLint: Generalizable Idiomatic Code Quality Analysis Through Instruction-Following and Easy-to-Hard Generalization

### v1

@@ A.3.1 Expanded Results on the PEP Hard Idiom Benchmark
MetaLint (SFT) 0.4333 0.0821 0.1381 0.0432 0.0221 0.0292
MetaLint (SFT+RS-DPO) 0.7031 0.7043 0.7037 0.3536 0.193 0.2497
MetaLint w CoT (RS-SFT) 0.7615 0.3689 0.497 0.2785 0.1437 0.1896
@@
o3-mini 0.8939 0.5845 0.7068 0.3169 0.2361 0.2706
o4-mini 0.9667 0.5943 0.7361 0.4131 0.3164 0.3584
Table 7: Result on the hard PEP benchmark to measure easy to hard generalization.
A.3.2 Statistical Significance of Results on the PEP Hard Idiom Benchmark
@@
A.3.3 Failure Analysis of MetaLint CoT Model VS Non CoT Model
@@
We also find similar issues in idioms like PEP 487, which discourages the use of metaclasses for simple customization tasks that could be handled via __init_subclass__ or __set_name__. The CoT model often misclassifies such “simple” use cases as complex.
@@
… the decorator is just a variable that’s a partial function. …
@@
is about using asynchronous generators with ”async for”
@@
don’t see any for loops that are using ’async for’. …
@@
complex … there’s a class called FirestoreClientMeta,
@@
with ”__init_subclass__” or
”__set_name__”)

### v2

@@ E.1 Expanded Results on the PEP Hard Idiom Benchmark
GPT-oss-20b 0.8377 0.3531 0.4968 0.251 0.1695 0.2024
GPT-oss-120b 0.9157 0.6456 0.7573 0.3991 0.3331 0.3631
Qwen3-4B MetaLint (SFT) (Ours) 0.4333 0.0821 0.1381 0.0432 0.0221 0.0292
Qwen3-4B MetaLint w CoT (RS-SFT) (Ours) 0.7615 0.3689 0.497 0.2785 0.1437 0.1896
Llama3.2-3B-Instruct MetaLint (SFT) (Ours) 0.5627 0.259 0.3547 0.1066 0.0509 0.0689
Llama3.2-3B-Instruct MetaLint (SFT+RS-DPO) (Ours) 0.6368 0.5614 0.5965 0.2364 0.1263 0.1647
o3-mini 0.8939 0.5845 0.7068 0.3169 0.2361 0.2706
o4-mini 0.9667 0.5943 0.7361 0.4131 0.3164 0.3584
@@
GPT-5 (high) 0.913 0.5673 0.6998 0.4397 0.4257 0.4326
Model Detection Localization
$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
Qwen3-4B 0.538 0.2637 0.3539 0.1396 0.1479 0.1436
Qwen3-4B + SFT 0.7686 0.3178 0.4497 0.2976 0.296 0.2968
Qwen3-4B + SFT + RS-DPO 0.7469 0.8315 0.7869 0.6527 0.6696 0.6611
Qwen3-4B w CoT 0.8812 0.6854 0.771 0.5049 0.4878 0.4962
Qwen3-4B w CoT + RS-SFT 0.935 0.8183 0.8727 0.6639 0.65 0.6569
Qwen3-4B w CoT + RS-SFT + RS-DPO 0.9234 0.8643 0.8929 0.771 0.7571 0.764
Llama3.2-3B-Instruct 0.5092 0.5286 0.5187 0.1371 0.3 0.1882
Llama3.2-3B-Instruct + SFT 0.6793 0.3598 0.4704 0.3424 0.3485 0.3454
Llama3.2-3B-Instruct + SFT + RS-DPO 0.654 0.6468 0.6504 0.491 0.4788 0.4848
Table 18: Cross-Idiom Generalization on Python Ruff Idioms: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B. Models are trained on easy synthetic Python Ruff idioms and tested on other Ruff idioms with varying levels of transfer (section 4.2). Best score across the compared training setups per model are bolded.
Model Detection Localization
$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
Qwen3-4B 0.5267 0.1715 0.2587 0.0954 0.0824 0.0884
Qwen3-4B + SFT 0.4333 0.0821 0.1381 0.0432 0.0221 0.0292
Qwen3-4B + SFT + RS-DPO 0.7031 0.7043 0.7037 0.3536 0.193 0.2497
Qwen3-4B w CoT 0.8154 0.3986 0.5354 0.2625 0.1467 0.1882
Qwen3-4B w CoT + RS-SFT 0.7615 0.3689 0.497 0.2785 0.1437 0.1896
Qwen3-4B w CoT + RS-SFT + RS-DPO 0.9303 0.4958 0.6468 0.3482 0.2169 0.2673
Llama3.2-3B-Instruct 0.7042 0.214 0.3283 0.0691 0.0798 0.0741
Llama3.2-3B-Instruct + SFT 0.5627 0.259 0.3547 0.1066 0.0509 0.0689
Llama3.2-3B-Instruct + SFT + RS-DPO 0.6368 0.5614 0.5965 0.2364 0.1263 0.1647
Table 19: Easy-to-Hard Generalization on PEP Idioms: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B. Models are trained on easy synthetic Python Ruff idioms and tested on hard manually curated PEP idiom detection data which can't be handled by linters or static analyzers (section 4.3). Best score across the compared training setups per model are bolded.
E.2 Statistical Significance of Results on the PEP Hard Idiom Benchmark
@@
E.3 Failure Analysis of MetaLint CoT Model VS Non CoT Model
@@
We also find similar issues in idioms like PEP 487, which discourages the use of metaclasses for simple customization tasks that could be handled via __init_subclass__ or __set_name__. The CoT model often misclassifies such ``simple'' use cases as complex.
@@
… the decorator is just a variable that's a partial function. …
@@
is about using asynchronous generators with "async for"
@@
don't see any for loops that are using 'async for'. …
@@
complex … there's a class called FirestoreClientMeta,
@@
with "__init_subclass__" or
"__set_name__")
@@
Model Transfer Detection Localization
$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Det}$ $R_{Det}$ $F_{Det}$
Llama3.2-3B-Instruct PMD $\rightarrow$ PMD 0.0457 0.0079 0.0134 0.0015 0.0022 0.0017
Llama3.2-3B-Instruct + SFT 0.2251 0.4421 0.2983 0.2822 0.2778 0.28
Llama3.2-3B-Instruct + SFT + RS-DPO 0.4395 0.8908 0.5886 0.593 0.5969 0.5949
Llama3.1-8B-Instruct PMD $\rightarrow$ PMD 0.3656 0.4015 0.3827 0.1253 0.131 0.1281
Llama3.1-8B-Instruct + SFT 0.2264 0.4508 0.3014 0.3201 0.3152 0.3177
Llama3.1-8B-Instruct + SFT + RS-DPO 0.4427 0.9191 0.5976 0.6506 0.6709 0.6606
Llama3.2-3B-Instruct PMD $\rightarrow$ JEP 0.3855 0.0096 0.0187 0.0005 0.0004 0.0005
Llama3.2-3B-Instruct + SFT 0.2286 0.4072 0.2928 0.1626 0.1336 0.1467
Llama3.2-3B-Instruct + SFT + RS-DPO 0.4903 0.8338 0.6175 0.4216 0.3333 0.3721
Llama3.1-8B-Instruct PMD $\rightarrow$ JEP 0 0 0 0 0 0
Llama3.1-8B-Instruct + SFT 0.2166 0.3724 0.2739 0.1455 0.1142 0.128
Llama3.1-8B-Instruct + SFT + RS-DPO 0.4964 0.8047 0.614 0.4615 0.3395 0.3912
Llama3.2-3B-Instruct JEP $\rightarrow$ JEP 0.3855 0.0096 0.0187 0.0005 0.0004 0.0005
Llama3.2-3B-Instruct + SFT 0.9567 0.8411 0.8952 0.7837 0.754 0.7686
Llama3.2-3B-Instruct + SFT + RS-DPO 0.9406 0.86 0.8985 0.7859 0.7651 0.7753
Llama3.1-8B-Instruct JEP $\rightarrow$ JEP 0 0 0 0 0 0
Llama3.1-8B-Instruct + SFT 0.9658 0.8466 0.9023 0.809 0.7844 0.7965
Llama3.1-8B-Instruct + SFT + RS-DPO 0.9308 0.8686 0.8986 0.8131 0.7756 0.7939
Llama3.2-3B-Instruct JEP $\rightarrow$ PMD 0.0457 0.0079 0.0134 0.0015 0.0022 0.0017
Llama3.2-3B-Instruct + SFT 0.3722 0.2708 0.3152 0.0574 0.0869 0.0692
Llama3.2-3B-Instruct + SFT + RS-DPO 0.4322 0.4054 0.4183 0.0878 0.1222 0.1022
Llama3.1-8B-Instruct JEP $\rightarrow$ PMD 0.3656 0.4015 0.3827 0.1253 0.131 0.1281
Llama3.1-8B-Instruct + SFT 0.3514 0.2229 0.2728 0.0383 0.0753 0.0508
Llama3.1-8B-Instruct + SFT + RS-DPO 0.436 0.4898 0.4613 0.0831 0.1351 0.1029
Table 28: Cross-Idiom Generalization on JEP & PMD Idioms: Effect of different MetaLint training setups (SFT and RS-DPO) on Llama3.2-3B-Instruct (Table 28). The transfer column indicates training and test data on the left and right side of the arrow. Best score across the compared training setups per model are bolded.